Hi, can you work out this equation? If so how many possible solutions are there? x^2 −3x+10=0
Thanks a lot! :)

Answer by oscargut(2103)

d = b^2-4ac = (-3)^2-4(1)(10) = 9-40 = -31 < 0
Then the equation has no real solutions


I can help you with more questions at: mthman@gmail.com
Thanks